A Professor of Economics from the Department of Economics Education of the Faculty of Social Sciences Education of the University of Education, Winneba, has estimated the demand and supply functions for students’ accommodation on campus to be respectively:
10P + 3Q = 60 and
P = Q – 0.5
Further, the professor claims that the average cost of Hostel operators is

If the hostel operators decide to maximize sales revenue instead of profit, show how this will affect consumer’s surplus. What is the producer’s surplus under pure competition at the equilibrium point?
1
Expert's answer
2021-08-29T16:51:08-0400

If the hostel operators decide to maximize sales revenue instead of profit, then this will increase the output and consumer’s surplus.

In equilibrium Qd = Qs and Pd = Ps, so:

10P + 3Q = 60, P = 6 - 0.3Q.

6 - 0.3Q = Q - 0.5,

Q = 5 units,

P = 5 - 0.5 = 4.5.

The producer’s surplus under pure competition at the equilibrium point is:

"PS = 0.5\u00d74.5\u00d7(5 + 0.5) = 12.375."